26 September 2014: Friday

LDAP:

  • One Issue in nightly update:
    • Former Student: did the dance.
  • CoM/IS retrieval stuff.

Backups:

  • Data Domain at 0.0 free again. 🙁
    • Compression has been reduced since we changed on 8/6 how we do the backups.
    • Defined a tape pool for the email backups and have shifted the backups to that (ugh!)
    • Opening issue with EMC/DataDomain about this…
    • Oh my goodness – there is 7 TB to be cleaned… but I thought I did that yesterday :/
    • EMC sees nothing wrong from the daily auto support logs – though they are doing an analysis of them and will get back to me.
    • Network Services reports the path is clean – no fragmentation, MTU 9000 the whole way.
    • Jim and I looked at things – nothing popped out – not major change in any settings on the data domain and the actual save set files look to be the same (good old tar format).
  • SR 65904996 (S:\ disk not being skipped):
    • proof that S:\ is not critical was given yesterday.
    • nsrget zip file was sent this morning.
    • No trouble found… please run “save -o VSS:LCV=yes” if it happens again!
  • Clones – last one finished yesterday.
  • Qualstar – tape drives turned over to FedEx.

Tags:

Leave a Reply

You must be logged in to post a comment.